OpenStudy (anonymous):

What is the best way to evaluate log base 2 ^16 ^23 without a calculator?

OpenStudy (anonymous):

\[\log_{2}16^{23} = 23\log_{2}16 \]

OpenStudy (anonymous):

\[16=2^4\] \[23\log_{2}2^4 = 23*4=92 \]